OpenStudy (anonymous):

2 log 50= 3 log 25+ log (x-2) (can anyone help)

OpenStudy (mathstudent55):

a log b = log (b^a) log a + log b = log (ab)

OpenStudy (anonymous):

i still dont understand can help me solve this problem

OpenStudy (mathstudent55):

apply the first rule above to the left side and to the first term of the right side log (50^2) = log (25^3) + log (x-2)

OpenStudy (anonymous):

log 2500 for 50^2

OpenStudy (mathstudent55):

Now apply the second rule to teh right side: log (50^2) = log [(25^3)(x - 2)] Now you have log a = log b, which means a = b, so: 50^2 = (25^3)(x - 2) and no more logs. Can you finish it now?

OpenStudy (anonymous):

so all i have to do is combine this problem together and ill have the answer 50^2 = (25^3)(x - 2)

OpenStudy (mathstudent55):

First divide both sides by 25^3 and solve for x

OpenStudy (anonymous):

625

OpenStudy (anonymous):

@mathstudent55 is it 625

OpenStudy (anonymous):

@Geckoh2o can you help with the next step

OpenStudy (mathstudent55):